OpenStudy (anonymous):

SOMEONE HELP ME D:!!! What value of x does f(x)=x^2-lnx have an absolute minimum?

OpenStudy (anonymous):

Take the derivative and then set the differential equation to zero. That should mark either a maximum or a minimum. What depends is the concavity of the original function.

OpenStudy (anonymous):

f'(x) = 2x - 1/x = 0 2x^2 - 1 = 0 x = +- sqrt(1/2)

OpenStudy (anonymous):

the negative value of x is inadmissable because a negative log is undefines soe the only turning point is at x = sqrt(1/2) second derivative is 2 + 1/x^2 which is postive for x= sqrt1/2 so this gives a minimum so x^2 - ln x has a minimum value when x = sqrt(1/2) if u go to wolframalpha.com and enter the fuction it will show you the graph and u'll see it is an absolute minimum